OpenStudy (anonymous):

I NEED HELP STAT Caleb and Sergio stacked boxes on a shelf. Caleb lifted 14 boxes and Sergio lifted 12 boxes. The boxes that Sergio lifted each weighed 10 lb more than the boxes Caleb lifted. Which expression represents the total number of pounds Sergio lifted? Let c represent the weight of each box Caleb lifted. A. 120c B. c + 10 C. 12(c + 10) D. 10(c − 12)

OpenStudy (anonymous):

Two stores sell the same television for the same original price. Store A advertises that the television is on sale for 30% off the original price. Store B advertises that it is reducing the television’s price by $250. When Allison compares the sale prices of the television in both stores, she concludes that the sale prices are equal. Which equation models this situation? Let p represent the television’s original price. A. 0.3p = p – 250 B. 0.7p = p – 250 C. 0.3p = p + 250 D. 0.7(p – 250) = p

OpenStudy (anonymous):

Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 4 h, and Car B traveled the distance in 4.5 h. Car B traveled 5 mph slower than Car A. How fast did Car B travel? mph

OpenStudy (anonymous):

Louann and Carla received equal scores on a test made up of multiple choice questions and an essay. Louann got 14 multiple choice questions correct and received 24 points for her essay. Carla got 16 multiple choice questions correct and received 14 points for her essay. How many points was each multiple choice question worth? points

OpenStudy (anonymous):

In May, Liam and Charlie had the same amount of money in their savings accounts. In June, Liam deposited $140 into his account. Charlie said he increased the money in his account by 5%. When they compared their balances, they found that they were still equal. How much money did they both have in their accounts in May? $

OpenStudy (anonymous):

@Capnkookcx

OpenStudy (♪chibiterasu):

We are not going to give you answers to your exam, it is Strictly against our Code of Conduct: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

ok @♪Chibiterasu

OpenStudy (anonymous):

can you atleast help me out

OpenStudy (anonymous):

@♪Chibiterasu

OpenStudy (anonymous):

@mary212129

OpenStudy (anonymous):

I am not going to give u answers either

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

i need help just walk me thourh

OpenStudy (jameshorton):

U NEED HELP